# Passage-traversing optimal path planning with sampling-based algorithms

### v1

@@ 1 Introduction
Motion and path planning makes a fundamental component in robotics and also many other disciplines. It resolves the basic problem of finding feasible paths connecting the start and goal states in a configuration space populated with infeasible portions, i.e., obstacles. Though path planning is hard from the algorithmic and computational perspectives, its prominent importance has driven the development of diverse path planning methods over the past decades, e.g., grid-based search, artificial potential fields, and sampling-based algorithms (LaValle, 2006; González et al., 2015; Paden et al., 2016; Orthey et al., 2023). Early studies mainly addressed path feasibility and only aimed at collision-free paths. Path optimality was not explicitly considered or simply referred to as path length. Up to date, the dominant planning paradigm has shifted to optimal path planning which seeks feasible paths optimizing specified objectives (Karaman and Frazzoli, 2011; Orthey et al., 2023).
@@
Despite the prevalence of optimal path planning, limited objective formalisms have been investigated. Path length, formally defined as the total variance of a path, is the most popular objective because it commonly reflects the time or other forms of costs to traverse the path. In practice, there are many other desired path properties such as good path smoothness and reasonable terrain selection (Rowe and Alexander, 2000). A fundamental one among these is to provide sufficient accessible free space along the entire path length (Sakcak and LaValle, 2021; Huang et al., 2023, 2024a). Accessible free space is reminiscent of the broadly used path clearance, the shortest distance between the path and obstacles, but represents a more general and comprehensive concept to gauge holistic accessible free space for a path. The requirement for it arises commonly in path planning in order to accommodate robots and their motions. For instance, manipulation in cluttered sites prefers large free space along the manipulated object’s path to reduce collision risk and manipulation complexity (Huang et al., 2023, 2024a). In navigation, large accessible free space along paths is required for the safety and agility of mobile robots (Mao et al., 2024; Wang et al., 2025). However, optimizing accessible free space in path planning remains under-investigated.
@@
To grant sufficient free space in path planning, clearance between robots and obstacles is embedded by inflating their volumes in collision check, e.g., (Lynch and Park, 2017; Paliwal and Kala, 2018). Alternatively, path clearance is formulated as a path cost term (Jaillet et al., 2010; Davoodi et al., 2015, 2020) or is enlarged by retracting planned paths (Geraerts and Overmars, 2007). Many works aim to find the maximum clearance, e.g., using visibility-Voronoi complex (Wein et al., 2007), corridors in Voronoi diagrams (Geraerts, 2010), and medial axes of the free space (Lien et al., 2003; Denny et al., 2014). Complex processing is required in geometric computations, costmap building, and path modifications in these methods. They also suffer from restrictive scalability in high-dimensional spaces, even in 3D space. More importantly, clearance cannot characterize the holistic accessible free space since it only reflects the most constrained situation. Maximizing path clearance alone does not ensure sufficient free space along the entire path.
@@
    A novel and general passage detection and associated free space decomposition method based on proximity graphs. The detected passages and cells are intuitive, sparse, and informative in characterizing the obstacle distribution and free space accessibility. This decomposition enables a fast passage-traversing check for paths.
@@
    Efficient algorithms for proposed methods with thorough analyses of probabilistic completeness, asymptotic optimality, and computational complexity. Passage and cell detection using Gabriel graphs is introduced for fast preprocessing. In PTOPP, the core module of rapid passage traversal check for paths is realized by positioning samples in subdivided cells, making PTOPP efficient.
@@
    Elaborate algorithm implementation and extensive experimental results for effectiveness and efficacy validation. For the common 2D and 3D workspace in robotics, all the proposals are implemented for thorough performance tests, which reveals the generality and applicability of PTOPP as a new paradigm of path planning.
A preliminary study called passage-aware optimal path planning was reported as a submodule in homotopic path set planning in (Huang et al., 2024a). This article investigates PTOPP independently and systematically, contextualizing it in general path planning problems. Substantial new theories, methods, and algorithms make PTOPP a self-contained new path planning paradigm.
@@ 2.1 Passages in path planning
While there is no unified definition of passages, they are generally recognized as passable regions in the configuration space whose removal changes the connectivity of the space (Hsu et al., 1998; Sun et al., 2005; Wang et al., 2018; Ruan et al., 2022). Specific characterizations of passages exist such as the notion of visibility sets (Barraquand et al., 1996; Hsu et al., 1997) and concatenation of a series of maximal spheres in free space (Guo et al., 2023). Frequently studied is narrow passage detection. The bridge detection tests based on the randomized bridge builder (RBB) (Hsu et al., 2003; Sun et al., 2005; Wang et al., 2010) and its variants, e.g., the randomized star builder (RSB) (Zhong and Su, 2013), bridge line tests (Lee et al., 2014), and the grid bridge builder (GBB) (Guo et al., 2023), are arguably the most prominent method. If the midpoint of two nearby in-collision points is collision-free, its neighborhood is marked as a passage region. Narrow passages are also identified as regions of large velocities in potential flows (Kazemi and Mehrandezh, 2004), high potentials in artificial potential fields (Aarno et al., 2004), and watersheds in decomposed space (Van den Berg and Overmars, 2005). A sample is considered in a narrow passage if the probabilistic roadmap (PRM) connected component it lies in has a small size (Shi et al., 2014). Using density-based spatial clustering, passage channels are detected as the sample clusters in narrow-passage-aware Gaussian sampling processes (Huang et al., 2024b).
@@
Narrow passages impose severe challenges for sampling-based path planning since the sampling probability sharply deteriorates inside them, rendering it hard to capture the connectivity of the sampled configuration space. Most works are contributed to sampling strategies for efficient path finding in the presence of narrow passages. In (Hsu et al., 1998), infeasible samples and connections of PRMs in the dilated free space are adjusted to the original free space via local resampling. Similarly, the free space is expanded by shrinking robots and obstacles in roadmap construction, and colliding portions are repaired by retraction in (Saha et al., 2005; Hsu et al., 2006). For sampling strategies, the Gaussian sampler is proposed in (Boor et al., 1999) in order to densify samples close to obstacles for better space coverage. Obstacle-based PRM (OBPRM) (Amato et al., 1998; Yeh et al., 2012) generates samples near obstacles for better connectivity in narrow passages. Another common avenue is to increase the sampling probability in detected narrow passages (Kazemi and Mehrandezh, 2004; Aarno et al., 2004; Van den Berg and Overmars, 2005) or to bias sampling to passages (Zucker et al., 2008). Exploiting obstacle or sample distribution information, the expansion methods of rapidly-exploring random tree (RRT) in different positions are designed in (Rodriguez et al., 2006; Tahirovic and Ferizbegovic, 2018) to move through narrow areas efficiently. Retracting invalid samples to obstacle boundaries (Zhang and Manocha, 2008) or passages (Lee et al., 2014) also helps to find paths traversing passages. In Triple RRTs (Wang et al., 2010; Zhong and Su, 2012), an RRT rooted at the midpoint in the bridge test attempts to connect to two RRTs rooted at the start and goal respectively in RRT-Connect. The analogous idea of growing RRTs in narrow passages is utilized within PRMs in (Shi et al., 2014). Explicit geometrical modeling of robots, environments (Ruan et al., 2022), and learning-based approaches for sampling and planning (Wang et al., 2018; Li and Dantam, 2023; Li et al., 2024) are also employed to address narrow passage problems.
@@ 2.2 Path clearance in path planning
Admitting sufficient accessible free space along the path is a fundamental requirement in robot path planning. This requirement is commonly characterized by path clearance, i.e., the shortest distance between the path and obstacles. A basic avenue to attaining clearance is by dilating obstacles and robots in collision check (Lynch and Park, 2017; Paliwal and Kala, 2018). The resulting path is automatically endowed with the prespecified clearance. Many works achieve this by constructing generalized Voronoi diagrams (medial axes of the free space) (Mark et al., 2008; Geraerts, 2010). Paths are searched on Voronoi diagrams to possess maximum clearance (Ó’Dúnlaing and Yap, 1985; Bhattacharya and Gavrilova, 2008). A hybrid diagram of the visibility graph and the Voronoi diagram enables the tradeoff between path length and clearance (Wein et al., 2007). Without explicitly computing medial axes, Medial Axis PRM (MAPRM) (Wilmarth et al., 1999; Lien et al., 2003) and Medial Axis RRT (MARRT) (Denny et al., 2014) adapt path nodes towards medial axes so that the formed roadmaps and trees have large clearance. Uniformly sampling on medial axes in PRM is presented in (Yeh et al., 2014). In (Guibas et al., 1999; Holleman and Kavraki, 2000), samples in PRM are restricted near the medial axes to efficiently sample narrow areas and increase path clearance. Similarly, skeleton-biased locally-seeded RRT (skilled-RRT) (Dong et al., 2018) leverages the geometric path on medial axes to bias RRT growth locally.
@@
Alternatively, clearance is optimized in path postprocessing locally performed without changing the path’s homotopy class. The external potential field is applied to push paths away from obstacles and maintain clearance in (Brock and Khatib, 2002). Path clearance gets increased by retracting path points to medial axes (Geraerts and Overmars, 2007) or introducing path points of larger clearance (Davoodi et al., 2015). Apart from explicitly specifying or adjusting path clearance, optimization problems are often formulated. Given roadmaps or grid maps, multi-objective graph-search, e.g., multi-objective A (Stewart and White III, 1991; Mandow and De la Cruz, 2005), bi-objective Dijkstra’s algorithm (Sedeño-Noda and Colebrook, 2019), and multi-objective evolutionary methods (Castillo et al., 2007; Ahmed and Deb, 2013), are exploited with a clearance objective. For continuous space, optimal RRT (RRT) and multi-objective evolutionary optimization are blended to asymptotically produce Pareto optimal solutions in (Yi et al., 2015). The weighted measure minimizing path length while maximizing path clearance is utilized in path search on Voronoi diagrams in (Wein et al., 2008). Pareto optimal solutions of the same bi-objective are found by a modified Dijkstra’s algorithm in (Davoodi et al., 2020). The integral of the reciprocal of clearance over the path length is minimized by sampling-based planners (Strub and Gammell, 2022). By introducing the Pareto optimal paths graph adopted from the visibility-Voronoi complex, the complete set of Pareto optimal solutions is obtainable in (Sakcak and LaValle, 2021), making the tradeoff between path length and clearance convenient to take.
@@ 2.3 Cell decomposition in path planning
Another relevant research topic is cell decomposition for path planning. It is a common practice to decompose the free space into a union of simple and non-overlapping subdivisions, i.e., cells, for better solvability in path planning problems (Brooks and Lozano-Perez, 1985). Grid search and sampling-based path planning methods are built on bitmap environment representation with uniform cells (LaValle et al., 2004; Tsardoulias et al., 2016) or regular cells of adaptive sizes (Chen et al., 1995; Van den Berg and Overmars, 2005; Lingelbach, 2004; Patle et al., 2019). Since the environment may not be precisely characterized by grids, e.g., irregular boundaries, grid decomposition is classified as approximate decomposition (Choset, 2001; Siegwart et al., 2011; Galceran and Carreras, 2013). More related to this work is exact cell decomposition that exploits environment features to obtain a complete environment representation. Classical decomposition approaches include triangulation and trapezoidal/vertical decomposition in 2D space as well as cylindrical decomposition in higher dimensional spaces (LaValle, 2006; Kloetzer and Ghita, 2011; Mark et al., 2008; Siciliano and Khatib, 2016; Aggarwal and Kumar, 2020). They utilize geometric features in polygonal obstacle regions, such as vertices, to construct cells. The adjacency graph encoding cell adjacency relationships is employed in graph-search path planning. Roadmaps are also built using nodes located in cells, e.g., centroids and boundary midpoints, as milestones for path planning. There exist other avenues like Morse decomposition (Choset and Pignon, 1998; Acar et al., 2002) and visibility-based decomposition (LaValle et al., 1997; Choset et al., 2005) based on changes of space connectivity and line of sight, which are complex to implement. A systematic comparison of different cell types, graph weights, and waypoint selections on cells in cell decomposing-based path planning can be found in (Kloetzer et al., 2015; Gonzalez et al., 2017).
@@
Path planning based on cell decomposition is commonly restrictive regarding path adjustability. In addition, most cell decomposition methods scale poorly in high-dimensional space, even 3D space. While partitioned cells make feasible paths easy to find and constraints, e.g., obstacle avoidance, easy to impose (Deits and Tedrake, 2015; Ryou et al., 2021), paths are not easily optimizable even for basic metrics such as the path length (Kloetzer et al., 2015). Cell decomposition generally gives rise to coarse paths returned by the shortest path search in grids or graphs. Integration with other optimal planners to unleash its potential in optimal path planning has been under-explored. Conventional usages of cells for roadmap construction (LaValle, 2006; Mark et al., 2008; Siciliano and Khatib, 2016; Aggarwal and Kumar, 2020) and direct integration in sampling-based path planning methods as samples (Rosell and Iniguez, 2005; Samaniego et al., 2017) suffer from low resolution unable to fulfill demanding planning requirements well. This article presents a novel and natural cell decomposition resulting from the narrow passage detection results, which is utilized in sampling-based optimal planners for fast sample positioning. Path’s passage traversal status is efficiently retrieved in PTOPP.
@@ 3 Passage detection and cell decomposition using Gabriel graphs
This section first presents a general PTOPP formulation as background. As essential preprocessing, an efficient passage detection and free space decomposition procedure based on proximity graphs and exploiting the inherent obstacle distribution properties is then introduced.
Figure 1: Example of dense passage distributions (orange segments) using the pure visibility condition in a laboratory for drone swarm flight, e.g., (Mao et al., 2024).

### v2

@@ 1 Introduction
Motion and path planning makes a fundamental component in robotics and also many other disciplines. It resolves the basic problem of finding feasible paths connecting the start and goal states in a configuration space populated with infeasible portions, i.e., obstacles. Though path planning is hard from the algorithmic and computational perspectives, its prominent importance has driven the development of diverse path planning methods over the past decades, e.g., grid-based search, artificial potential fields, and sampling-based algorithms (lavalle2006planning; gonzalez2015review; paden2016survey; orthey2023sampling). Early studies mainly addressed path feasibility and only aimed at collision-free paths. Path optimality was not explicitly considered or simply referred to as path length. Up to date, the dominant planning paradigm has shifted to optimal path planning which seeks feasible paths optimizing specified objectives (karaman2011sampling; orthey2023sampling).
@@
Despite the prevalence of optimal path planning, limited objectives have been investigated. Path length, formally defined as the total variance of a path, is the most popular objective because it commonly reflects the time or cost to traverse the path. In practice, there are many other desired path properties such as good path smoothness and reasonable terrain selection (rowe2000finding). A fundamental one among these is to provide sufficient accessible free space along the entire path length (sakcak2021complete; huang2023deformable; huang2024homotopic). Accessible free space is reminiscent of the broadly used path clearance, the shortest distance between the path and obstacles, but represents a more general and comprehensive concept to gauge holistic accessible free space for a path. The requirement for it arises commonly in path planning in order to accommodate robots and their motions. For instance, manipulation in cluttered sites prefers large free space along the manipulated object’s path to reduce collision risk and manipulation complexity (huang2023deformable; huang2024homotopic). In navigation, large accessible free space along paths is required for the safety and agility of mobile robots (mao2024optimal; wang2025fast). However, optimizing accessible free space in path planning remains under-investigated.
@@
To grant sufficient free space in path planning, clearance between robots and obstacles is embedded by inflating their volumes in collision check, e.g., (lynch2017modern; paliwal2018maximum). Alternatively, path clearance is formulated as a path cost term (jaillet2010sampling; davoodi2015clear; davoodi2020path) or gets enlarged by retracting planned paths (geraerts2007creating). Many works aim to find the maximum clearance, e.g., using visibility-Voronoi complex (wein2007visibility), corridors in Voronoi diagrams (geraerts2010planning), and medial axes of the free space (lien2003general; denny2014marrt). Complex processing is required in geometric computations, costmap building, and path adjustments in these methods. They also suffer from restrictive scalability in high-dimensional spaces, even in 3D space. More importantly, clearance cannot characterize the holistic accessible free space since it only reflects the most constrained situation. Maximizing clearance alone does not ensure adequate free space along the entire path.
@@
    A novel and general passage detection and associated free space decomposition method based on proximity graphs. The detected passages and cells are intuitive, sparse, and informative in characterizing the obstacle distribution and free space accessibility. This decomposition enables fast passage-traversing check for paths.
@@
    Efficient algorithms for PTOPP. Fast passage and cell detection using Gabriel graphs is introduced. In PTOPP, the core module of rapid passage traversal check for paths is realized by positioning samples in subdivided cells, making PTOPP efficient.
@@
    Elaborate algorithm implementation and extensive experimental results for effectiveness and efficacy validation. For the common 2D and 3D workspace, all the proposals are implemented for thorough performance tests, which reveals the generality and applicability of PTOPP as a new paradigm of path planning.
A preliminary study called passage-aware optimal path planning was reported as a submodule in homotopic path set planning in (huang2024homotopic). This article investigates PTOPP independently and systematically, contextualizing it in general path planning problems. Substantial new theories, methods, and algorithms make PTOPP a self-contained new path planning paradigm.
@@ 2.1 Passages in path planning
While there is no unified definition of passages, they are generally recognized as passable regions in the configuration space whose removal changes the connectivity of the space (hsu1998finding; sun2005narrow; wang2018learning; ruan2022efficient). Specific characterizations of passages exist such as the notion of visibility sets (barraquand1996random; hsu1997path) and concatenation of a series of maximal spheres in free space (guo2023obstacle). Frequently studied is narrow passage detection. The bridge detection tests based on the randomized bridge builder (RBB) (hsu2003bridge; sun2005narrow; wang2010triple) and its variants, e.g., the randomized star builder (RSB) (zhong2013robot), bridge line tests (lee2014selective), and the grid bridge builder (GBB) (guo2023obstacle), are arguably the most prominent method. If the midpoint of two nearby in-collision points is collision-free, its neighborhood is marked as a passage region. Narrow passages are also identified as regions of large velocities in potential flows (kazemi2004robotic), high potentials in artificial potential fields (aarno2004artificial), and watersheds in decomposed space (van2005using). A sample is considered in a narrow passage if the probabilistic roadmap (PRM) connected component it lies in has a small size (shi2014spark). Using density-based spatial clustering, passage channels are detected as the sample clusters in narrow-passage-aware Gaussian sampling processes (huang2024robot).
@@
Narrow passages impose severe challenges for sampling-based path planning since the sampling probability sharply deteriorates inside them, rendering it hard to capture the connectivity of the sampled configuration space. Most works are contributed to sampling strategies for efficient path finding in the presence of narrow passages. In (hsu1998finding), infeasible samples and connections of PRMs in the dilated free space are adjusted to the original free space via local resampling. Similarly, the free space is expanded by shrinking robots and obstacles in roadmap construction, and colliding portions are repaired by retraction in (saha2005finding; hsu2006multi). For sampling strategies, the Gaussian sampler is proposed in (boor1999gaussian) in order to densify samples close to obstacles for better space coverage. Obstacle-based PRM (OBPRM) (amato1998obprm; yeh2012uobprm) generates samples near obstacles for better connectivity in narrow passages. Another common avenue is to increase the sampling probability in detected narrow passages (kazemi2004robotic; aarno2004artificial; van2005using) or to bias sampling to passages (zucker2008adaptive). Exploiting obstacle or sample distribution information, the expansion methods of rapidly-exploring random tree (RRT) in different positions are designed in (rodriguez2006obstacle; tahirovic2018rapidly) to move through narrow areas efficiently. Retracting invalid samples to obstacle boundaries (zhang2008efficient) or passages (lee2014selective) also helps to find paths traversing passages. In Triple RRTs (wang2010triple; zhong2012triple), an RRT rooted at the midpoint in the bridge test attempts to connect to two RRTs rooted at the start and goal respectively in RRT-Connect. The analogous idea of growing RRTs in narrow passages is utilized within PRMs in (shi2014spark). Explicit geometrical modeling of robots, environments (ruan2022efficient), and learning-based approaches for sampling and planning (wang2018learning; li2023sample; li2024sampling) are also employed to address narrow passage problems.
@@ 2.2 Path clearance in path planning
Admitting sufficient accessible free space along the path is a fundamental requirement in robot path planning. This requirement is commonly characterized by path clearance, i.e., the shortest distance between the path and obstacles. A basic avenue to attaining clearance is by dilating obstacles and robots in collision check (lynch2017modern; paliwal2018maximum). The resulting path is automatically endowed with the prespecified clearance. Many works achieve this by constructing generalized Voronoi diagrams (medial axes of the free space) (mark2008computational; geraerts2010planning). Paths are searched on Voronoi diagrams to possess maximum clearance (o1985retraction; bhattacharya2008roadmap). A hybrid diagram of the visibility graph and the Voronoi diagram enables the tradeoff between path length and clearance (wein2007visibility). Without explicitly computing medial axes, Medial Axis PRM (MAPRM) (wilmarth1999maprm; lien2003general) and Medial Axis RRT (MARRT) (denny2014marrt) adapt path nodes towards medial axes so that the formed roadmaps and trees have large clearance. Uniformly sampling on medial axes in PRM is presented in (yeh2014umaprm). In (guibas1999probabilistic; holleman2000framework), samples in PRM are restricted near the medial axes to efficiently sample narrow areas and increase path clearance. Similarly, skeleton-biased locally-seeded RRT (skilled-RRT) (dong2018faster) leverages the geometric path on medial axes to bias RRT growth locally.
@@
Alternatively, clearance is optimized in path postprocessing locally performed without changing the path’s homotopy class. The external potential field is applied to push paths away from obstacles and maintain clearance in (brock2002elastic). Path clearance gets increased by retracting path points to medial axes (geraerts2007creating) or introducing path points of larger clearance (davoodi2015clear). Apart from explicitly specifying or adjusting path clearance, optimization problems are often formulated. Given roadmaps or grid maps, multi-objective graph-search, e.g., multi-objective A (stewart1991multiobjective; mandow2005new), bi-objective Dijkstra’s algorithm (sedeno2019biobjective), and multi-objective evolutionary methods (castillo2007multiple; ahmed2013multi), are exploited with a clearance objective. For continuous space, optimal RRT (RRT) and multi-objective evolutionary optimization are blended to asymptotically produce Pareto optimal solutions in (yi2015morrf). The weighted measure minimizing path length while maximizing path clearance is utilized in path search on Voronoi diagrams in (wein2008planning). Pareto optimal solutions of the same bi-objective are found by a modified Dijkstra’s algorithm in (davoodi2020path). The integral of the reciprocal of clearance over the path length is minimized by sampling-based planners (strub2022adaptively). By introducing the Pareto optimal paths graph adopted from the visibility-Voronoi complex, the complete set of Pareto optimal solutions is obtainable in (sakcak2021complete), making the tradeoff between path length and clearance convenient to take.
@@ 2.3 Cell decomposition in path planning
Another relevant research topic is cell decomposition for path planning. It is a common practice to decompose the free space into a union of simple and non-overlapping subdivisions, i.e., cells, for better solvability in path planning problems (brooks1985subdivision). Grid search and sampling-based path planning methods are built on bitmap environment representation with uniform cells (lavalle2004relationship; tsardoulias2016review) or regular cells of adaptive sizes (chen1995planning; van2005using; lingelbach2004path; patle2019review). Since the environment may not be precisely characterized by grids, e.g., irregular boundaries, grid decomposition is classified as approximate decomposition (choset2001coverage; siegwart2011introduction; galceran2013survey). More related to this work is exact cell decomposition that exploits environment features to obtain a complete environment representation. Classical decomposition approaches include triangulation and trapezoidal/vertical decomposition in 2D space as well as cylindrical decomposition in higher dimensional spaces (lavalle2006planning; kloetzer2011software; mark2008computational; siciliano2016springer; aggarwal2020path). They utilize geometric features in polygonal obstacle regions, such as vertices, to construct cells. The adjacency graph encoding cell adjacency relationships is employed in graph-search path planning. Roadmaps are also built using nodes located in cells, e.g., centroids and boundary midpoints, as milestones for path planning. There exist other avenues like Morse decomposition (choset1998coverage; acar2002morse) and visibility-based decomposition (lavalle1997finding; choset2005principles) based on changes of space connectivity and line of sight, which are complex to implement. A systematic comparison of different cell types, graph weights, and waypoint selections on cells in cell decomposing-based path planning can be found in (kloetzer2015optimizing; gonzalez2017comparative).
@@
Path planning based on cell decomposition is commonly restrictive regarding path adjustability. In addition, most cell decomposition methods scale poorly in high-dimensional space, even 3D space. While partitioned cells make feasible paths easy to find and constraints, e.g., obstacle avoidance, easy to impose (deits2015efficient; ryou2021multi), paths are not easily optimizable even for basic metrics such as the path length (kloetzer2015optimizing). Cell decomposition generally gives rise to coarse paths returned by the shortest path search in grids or graphs. Integration with other optimal planners to unleash its potential in optimal path planning has been under-explored. Conventional usages of cells for roadmap construction (lavalle2006planning; mark2008computational; siciliano2016springer; aggarwal2020path) and direct integration in sampling-based path planning methods as samples (rosell2005path; samaniego2017uav) suffer from low resolution unable to fulfill demanding planning requirements well. This article presents a novel and natural cell decomposition resulting from the narrow passage detection results, which is utilized in sampling-based optimal planners for fast sample positioning. Path’s passage traversal status is efficiently retrieved in PTOPP.
@@ 3 Passage detection and cell decomposition using Gabriel graphs
This section first presents a general PTOPP formulation as background. As essential preprocessing, an efficient passage detection and free space decomposition procedure based on proximity graphs of obstacles is then introduced.